A person invests 6000 dollars in a bank. The bank pays 6% interest compounded daily. To the nearest tenth of a year, how long must the person leave the money in the bank until it reaches 11100 dollars?
Answer by lwsshak3(11628)

A person invests 6000 dollars in a bank. The bank pays 6% interest compounded daily. To the nearest tenth of a year, how long must the person leave the money in the bank until it reaches 11100 dollars?
***
A=P(1+r/n)^nt, P=initial investment, r=interest rate, n=number of compounding periods per year
A=amt after t-years
..
assume 365 days/yr
11100=6000(1+.06/365)^365t
11100/6000=(1+.0001644)^365t
111/60=(1.0001644)^365t
take log of both sides
log(111/60)=365tlog(1.0001544)
t=log(111/60)/365log(1.0001544)=14.2
how long must the person leave the money in the bank? about 14 years